OpenStudy (anonymous):

The following polynomial represents a profit function for a certain production line, where x is a number of produced units, find a. the zeros and the multiplicity of each b. where the graph crosses or touches the x-axis c. number of turning points d. explain the meaning of the above, if any f (x) = x^2 (x + 2)^3 (x^2 - 1)

OpenStudy (anonymous):

\[f (x) = x^2 (x + 2)^3 (x^2 - 1)=f (x) = x^2 (x + 2)^3 (x+1)(x-1)\] zeros are displayed for you \[x^2=0\iff x=0\] the mutiplicity is 2 because of the exponent \[(x+2)^2=0\iff x=-2\] and here the multiplicy is 3 \[x+1=0\iff x=-1\] \[x-1=0\iff x=1\] both have multiplicty 1

OpenStudy (anonymous):

oops typo \[(x+2)^3=0\iff x=-2\] multiplicty is 3

OpenStudy (anonymous):

Thanks a bunch satellite 73

OpenStudy (anonymous):

By chance are you able to determine the number of turning point